# Collision Energy Dependence of Particle Ratios and Freeze-out Parameters in Ultra Relativistic Nucleus Nucleus Collisions

### v1

@@ III.1 LIKE MASS ANTIHYPERON TO HYPERON RATIOS
Experimental data show that all antibaryon-to-baryon ratios increase with $\sqrt{s_{NN}}$ and approach unity as we move towards the LHC energies, where the temperature $(T)$ almost saturates at 170 MeV and the chemical potential $(\mu_{B}$) nearly vanishes, indicating an almost baryon-antibaryon symmetric matter. At lower energies (i.e., SPS and AGS), the system maintains a large excess of baryons over antibaryons, resulting in all antibaryon-to-baryon ratios being significantly lower than unity Grebieszkow2004 . In Figs. 4-4, we have shown the energy dependence of antihyperon to hyperon ratios, i.e., $\frac{\bar{\mathrm{\Lambda}}}{\Lambda}$, $\frac{\bar{\mathrm{\Xi}}}{\mathrm{\Xi}}$, and $\frac{\bar{\mathrm{\Omega}}}{\mathrm{\Omega}}$. The minimum $\chi^{2}/\text{dof}$ values for the above three cases are 6.19, 0.81, and 3.43, respectively. The large value for the $\frac{\bar{\mathrm{\Lambda}}}{\Lambda}$ case is apparently due to very small error bars for the lower energy data points.
Figure 2:
Figure 3:
Figure 4:
Figs. 4-4 Variation of Λ¯/$\bar{\Lambda}/\Lambda$, Ξ¯/$\bar{\Xi}/\Xi$ and Ω¯/$\bar{\Omega}/\Omega$ with $\sqrt{s_{NN}}$
@@
The increasing antibaryon to baryon ratios with $\sqrt{s_{NN}}$ also suggests a higher degree of nuclear transparency in the URNNC at higher collision energies, where the bulk of the secondary matter is formed between the two receding nuclei due to high excitation of vacuum and hence is almost baryon symmetric. This effect becomes more prominent as we move towards the highest RHIC and further up to the LHC energies. In lower energy collisions, the bulk secondary matter contains a large contribution from the participant nucleons due to significant baryon stopping in these collisions. This observation is also supported by the decreasing value of net-protons ($p-\bar{p}$) measured in the experiments with increasing collision energy. The matter formed at lower RHIC energies indicates a situation where partial transparency exists in the nuclear collisions. This observation can provide insights into the evolving collision dynamics with increasing energy Abelev2010 . The curves obtained by our VDW-HRG model calculations exhibit good agreement with experimental data Afanasiev2002 ; Antinori2004 ; Appelshauser1998 ; Alt2005Omega ; Mischke2002 ; Aggarwal2011S ; Abbas2013b ; Abelev2010 . The energy-dependent behavior of antihyperon/hyperon ratios, depicted in the Figs. 4, 4, and 4 reveal a mass-hierarchical trend, with a saturation value of 1 being reached earlier for more massive hyperon species. This phenomenon underscores a preference for almost symmetric antihyperons-hyperons production in the case of multistrange hyperons, where the former, i.e., the $\Lambda$, contains only two light valence quarks, while the $\Xi$ contains only one light valence quark, the likely remnants from the initial nucleons. In the case of $\mathrm{\Omega}$, all (strange) valence quarks are newly generated Andronic2006 ; Bashir2015 . The generally unsatisfactory quality of thermal fits to the multi-strange baryons data at SPS energies cannot solely be attributed to these cases, as we shall see below Andronic2006 . These ratios (including $K^{-}/K^{+}$) increase sharply with respect to $\sqrt{s_{NN}}$ and then almost saturate at higher energies (above 200 GeV), reaching a value equal to 1.0 at LHC energy. This behavior shows that the production rates of anti-particles relative to particles continuously increase with increasing $\sqrt{s_{NN}}$ and become almost equal at LHC energy. The excellent agreement between our VDW-HRG model results and the experimental data validates the ability of the model to explain these particle ratios.
III.2 MESONIC RATIOS
Extending our analysis further to obtain the minimum $\chi^{2}/\text{dof}$ fit to the $K^{-}/K^{+}$ ratio data against $\sqrt{s_{NN}}$, we find that it can be best fitted only with a modified (lesser) value of the chemical freeze-out temperature where the parameter $e$ turns out to be 162 MeV instead of 172 MeV (obtained for baryons) while the remaining parameters remain unchanged. This seems to arise due to the later freeze-out of Kaons (including pions as we will see later) than baryons when the system has further cooled down due to expansion. The baryons therefore appear to freeze-out slightly earlier than the kaons due to their larger mean free path in the system. The best fit to the $K^{-}/K^{+}$ data gives $\chi^{2}/\text{dof}=1.28$. This clearly indicates the existence of two chemical freeze-out stages in the system, where the second freeze-out corresponds to the situation where mesons ($K^{-}$, $K^{+}$ and $\pi$) decouple from the system. In some other analyses, it has been suggested that using two chemical freeze-out temperatures improves the thermal model accuracy thereby enhancing its predictive power and better matching with experimental data Assis2013 ; Duarte2012 . Thus, the freeze-out temperature is an effective thermodynamic quantity linked to the system’s expansion.
The $\sqrt{s_{NN}}$ dependent theoretical curve for the $K^{-}/K^{+}$ ratio obtained using the VDW-HRG model calculations is shown in Fig 5 which is in good agreement with the experimental data Abelev2010 ; Abelev2009 ; Acharya2020 ; Adams2002S ; Abelev2013 ; Mischke2002 . The decline in the $K^{-}/K^{+}$ ratio as collision energy decreases is attributed to a rise in net baryon density, favoring the production of $K^{+}$ (having $u$ and $\bar{s}$ valence quarks) over $K^{-}$ (having $\bar{u}$ and $s$ valence quarks).
Figure 5: $\sqrt{s_{NN}}$ dependence of $K^{-}/K^{+}$
@@ III.2 MESONIC RATIOS
As the collision energy increases, the $K^{-}/K^{+}$ ratio exhibits a systematic rise, indicative of the pair production processes for $K^{+}$ and $K^{-}$ at higher energies in almost equal proportion through the reaction channels involving baryons and antibaryons in an almost baryon symmetric matter (having $\mu_{B}\sim 0$). The associated production mechanism yields $K^{+}$ through reactions like $\pi+N\rightarrow Y+K^{+}$ and the production of $K^{-}$ takes place through $\pi+\bar{N}\rightarrow\bar{Y}+K^{-}$, where $N$ represents a nucleon and $Y$ denotes a singly strange hyperon. On the other hand, reactions like $N+\bar{N}\rightarrow K^{-}+K^{+}$ and $\pi^{+}+\pi^{-}\rightarrow K^{-}+K^{+}$ will produce $K^{+}$ and $K^{-}$ in equal proportion Das2014 .
Despite the success of the HRG Model in describing many particle ratios, some unresolved issues still remain while describing the yields of the strange (antistrange) particles relative to the non-strange ones. For nearly a decade, the $K/\pi$, $\Lambda/\pi^{-}$, etc., ratios were not accurately described Beutler2008 ; Tawfik2016 . Pions being the lightest hadrons are produced copiously and carry the bulk of the entropy of the system. The dependence of the $K/\pi$ on $\sqrt{s_{NN}}$ has always been a subject of interest as it is a crucial indicator of the strangeness content of the system relative to its entropy in heavy-ion collisions and has implications for the quark-hadron phase transition. The observed increase and subsequent saturation of the $K/\pi$ ratio with increasing beam energies align with established trends and theoretical expectations Andronic2006 . Assuming simultaneous freeze-out for kaons and pions in our calculation, we have obtained our best fit to the experimental data on $K^{-}/\pi^{-}$ ratio dependence on $\sqrt{s_{NN}}$. The best-fitted curve obtained is shown in Fig 6 in blue color. It is obvious that the theoretical curve provides an unsatisfactory fit to the experimental dataAdams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 . However, such large disagreement has been reported earlier due to strangeness imbalance Castorina2016 ; Munzinger2003 ; Becattini2005 ; Bugaev2016 .
@@
The theoretical curve is seen to under-predict the experimental data at lower collision energies ($\sqrt{s_{NN}}<10\,\text{GeV}$) and over-predict it at higher energies. We thus introduce a strangeness imbalance factor $\xi_{s}$ which is defined at different energies (say for the $i^{th}$) as followsTawfik2021H ; Biswas2020 :
@@
Further, in order to make it convenient to apply this factor to other particle ratios as well for any given collision energy, we have used a parameterization method to represent the energy dependence of $\xi_{s}$. We have used the following ansatz to best represent the obtained $\sqrt{s_{NN}}$ dependent values of $\xi_{s}$:
@@
The best spline fit provides $\alpha=0.49$, $\beta=8.37\times 10^{6}$, $\gamma=0.26\,\text{GeV}$, and $\delta=4.63$.
Figure 7: $\sqrt{s_{NN}}$ dependence of $\xi_{s}$
In Fig 7, we have shown the energy dependence of $\xi_{s}$. Utilizing the ansatz of Eq. 39, we have redefined the strangeness imbalance corrected strange (antistrange) particle number densities in our calculation at different collision energies as
@@
The quantity $h$ represents the strangeness content of a given particle. For singly strange particles (i.e., kaons, antikaons, lambda, antilambda, sigma, antisigma) we have $h=1$; for doubly strange particles (i.e., cascades, anticascades) we have $h=2$; and similarly for triply strange particles (i.e., omega, antiomega) we have $h=3$. In Fig 6, the corrected theoretical curve for the $K^{-}/\pi^{-}$ ratio using the above ansatz in Eq. 39 is shown by the red curve, which has a very small $\chi^{2}/\text{dof}=0.04$, resulting from the correction procedure. Moreover, using the values of $\xi_{s}$ obtained from Eq. 39, we have also calculated the strangeness imbalance corrected $K^{+}/\pi^{+}$ ratio for different collision energies. In Fig 8, the theoretical curve for the $K^{+}/\pi^{+}$ ratio is shown in red, giving a reasonable value of $\chi^{2}/\text{dof}=0.16$. Interestingly, the theoretical curve also explains the horn structure at lower energies, $\sqrt{s_{NN}}\sim 7-8\,\text{GeV}$. The steep variation of the $(K^{+}/\pi^{+})$ and $(K^{-}/\pi^{-})$ ratios at lower energies reflects the threshold for strangeness production affected by the rapid initial increase in the temperature with $\sqrt{s_{NN}}$. A steep drop in the $K^{+}/\pi^{+}$ near the horn may arise due to the rapidly falling baryon chemical potential, as baryon-rich matter favors $K^{+}$ production, as discussed above Adams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 ; RecentRF ; NA61SHINE2020 ; Ahle1999P .
Figure 8: Variation of $K^{+}/\pi^{+}$ with $\sqrt{s_{NN}}$
Figure 9: Dependence of $\pi^{-}/\pi^{+}$ and $K^{-}/K^{+}$ on $\sqrt{s_{NN}}$
In Fig. 9, we have shown the energy dependence of the $\pi^{-}/\pi^{+}$ ratio. The green curve shows the theoretical fit, which matches the experimental data quite well with $\chi^{2}/\text{dof}=0.37$. It may be recalled here that in the calculation of all particle ratios, we have also applied the criteria of conserved charge to baryon ratio in the experiments in order to fix the electric chemical potential $\mu_{Q}$. We find that the $\pi^{-}/\pi^{+}$ ratio decreases with $\sqrt{s_{NN}}$ gradually and saturates towards unity as the system tends to become almost charge symmetric. In the same figure, we have again shown the $K^{-}/K^{+}$ ratio for comparison, as both ratios become equal ($\sim 1$) for very large $\sqrt{s_{NN}}$. At low beam energies, the $\pi^{-}/\pi^{+}$ ratios exhibit values larger than unity, possibly due to significant contributions to the positive charges from protons and other resonances, particularly the $\Delta$ baryons. For the ease of comparison we have also shown the $K^{-}/K^{+}$ ratio in the same figure. The distribution characteristics of the $\pi^{-}/\pi^{+}$ and $K^{-}/K^{+}$ ratios differ from each other however both ratios are seen approaching 1 due to nuclear transparency effects. The theoretical results are in reasonable agreement with the experimental data Bratkovskaya2004 ; Adams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 . This also explains the observation that the $K^{-}/\pi^{-}$ and $K^{+}/\pi^{+}$ ratios in Figs. 6 and 8 approach the same value ($\sim 0.15$) towards the highest RHIC and LHC energies.
In Figs. 10, 11, 12, 13, and 14, we have shown the results of our calculations along with the experimental data for several particle ratios, i.e., $\frac{\Lambda(\bar{\Lambda})}{\pi^{-}}$, $\frac{\Xi(\bar{\Xi})}{\pi^{-}(\pi^{+})}$, $\Omega/\pi$, $\Phi/\pi^{-}$, and $\Phi/K^{-}$. We find that a reasonable fit is obtained in all these cases, thus highlighting the need for consideration of the strangeness imbalance effect in the system. Corresponding experimental data are taken from the references Adam2020 ; Abelev2010 ; Adamczyk2017B ; Antii2003 ; RecentRF ; Alt2008 ; Aggarwal2011S ; Friese2004 ; E9172003 . The $\mathrm{\Phi}\left(s\bar{s}\right)$ mesons with hidden strangeness, having relatively small cross sections with other hadrons, are assumed to freeze-out earlier than Kaons and pions (i.e., along with baryons). Our model calculations are able to describe the $\phi/\pi$ and $\Omega/\pi$ ratios reasonably well. Our results are in better agreement than an earlier result Tiwari2013 . Due to the strangeness imbalance effect the like-mass anti-hyperon to hyperon as well as antikaon to kaon ratios are not altered, as their abundances in the system are equally affected. In their theoretical calculations the correction factor $\xi_{s}$ cancels.
Figure 10:
Figure 11:
Figure 12:
Figure 13:
Figure 14:
Figs. 10-14 Variation of $\Lambda(\bar{\Lambda})/\pi^{-}$, $\Xi/\pi$ , $\Omega/\pi^{-}$, $\phi/\pi^{-}$ and $\phi/K^{-}$ with $\sqrt{s_{NN}}$
As the protons (antiprotons) are the lightest baryons (antibaryons) hence most high-mass baryons (antibaryons) decay into protons (antiprotons). The $p/\pi$ ratio, therefore, characterizes well the overall baryon (antibaryon) production relative to the total particle multiplicity. To understand this, we have shown the $p/\pi^{+}$ and $\bar{p}/\pi^{-}$ ratios energy dependence in Fig 15. The theoretical curve appears to fit the data quite well with $\chi^{2}/\text{dof}=1.60$ and $4.20$, respectively. The larger value for the second case arises due to very small error bars in the experimental data Nandi2020 ; Aggarwal2011S ; Adamczyk2017B ; Abelev2010 .
Figure 15: Variation of $p/\pi$ with $\sqrt{s_{NN}}$
@@
We have analyzed the correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios in URNNC over a wide energy range from 5 GeV to 2.76 TeV. The correlation is best described by a power-law type relationship as $K^{-}/K^{+}=\left(\bar{p}/p\right)^{\nu}$. This can serve as validation for any thermal model as it indicates how kaon production relates to net baryon density. In Fig 16, we have shown the experimental data points along with curves obtained from different approaches. The blue curve shows the result of our calculations. The pink curve is obtained by considering the light valence quark compositions, thus providing $\nu\approx 0.33$ but fits the data somewhat poorly. For comparison, if we represent the theoretical blue curve by a power-law fit, this provides $\nu\approx 0.18$, which contrasts sharply with the light valence quark composition approach. However, if we obtain a direct fit to the data using the power law, we get $\nu\approx 0.21$, which is quite close to our theoretical results. This correlation is therefore well described by the VDW-HRG type approach in the framework of the grand canonical ensemble-based thermal model. Experimental data at different energies are taken from the references Bearden2003 ; Stiles2006 ; Das2014 ; Abelev2009 .
III.3 ENERGY DEPENDENCE OF THERMAL FREEZE-OUT PARAMETERS ($T$, $\mu_{B}$, $\mu_{s}$)

### v2

@@ IV.1 LIKE MASS ANTIBARYON TO BARYON RATIOS
Experimental data show that all antibaryon-to-baryon ratios increase with $\sqrt{s_{NN}}$ and approach unity as we move towards the LHC energies, where the temperature $(T)$ almost saturates at 164 MeV and the chemical potential $(\mu_{B}$) nearly vanishes, indicating an almost baryon-antibaryon symmetric matter. At lower energies (i.e., SPS and AGS), the system maintains a large excess of baryons over antibaryons, resulting in all the antibaryon-to-baryon ratios being significantly lower than unity Grebieszkow2004 . In Fig. 1, we have shown the energy dependence of antibaryon to baryon ratios, i.e., $\bar{p}/p$, Λ¯/$\bar{\Lambda}/\Lambda$, Ξ¯/$\bar{\Xi}/\Xi$, and Ω¯/$\bar{\Omega}/\Omega$. The minimum $\chi^{2}/\text{{dof}}$ values for the above four cases are 0.54, 0.79, 3.92, and 0.68, respectively. A relatively larger value in case of Ξ¯/$\bar{\Xi}/\Xi$ is due to the small error bar associated with the data point at 2.76 TeV.
Figure 1: Collision energy dependence of $\bar{p}/p$, Λ¯/$\bar{\Lambda}/\Lambda$, Ξ¯/$\bar{\Xi}/\Xi$, Ω¯/$\bar{\Omega}/\Omega$.
@@
The increasing antibaryon to baryon ratios with $\sqrt{s_{NN}}$ also suggests a higher degree of nuclear transparency in the URNNC at higher collision energies, where the bulk of the secondary matter is formed between the two receding nuclei due to high excitation of vacuum and hence is almost baryon symmetric. This effect becomes more prominent as we move towards the highest RHIC and further up to the LHC energies. In lower energy collisions, the bulk secondary matter contains a large contribution from the participant nucleons due to significant baryon stopping in these collisions. This observation is also supported by the decreasing value of net-protons ($p-\bar{p}$) measured in the experiments with increasing collision energy. The matter formed at lower RHIC energies indicates a situation where partial transparency exists in the nuclear collisions. This observation can provide insights into the evolving collision dynamics with increasing energy Abelev2010 . The curves obtained by our VDW-HRG model calculations exhibit good agreement with experimental data taken from various available sources Adamczyk2017B ; Ahle1999P ; Adams2002S ; Aggarwal2011S ; Aamodt2010M ; Afanasiev2002 ; Antinori2004 ; Appelshauser1998 ; Alt2005Omega ; Mischke2002 ; Abbas2013b ; Abelev2010 . The energy-dependent behavior of antihyperon/hyperon ratios, depicted in the Fig. 1 reveals a mass-hierarchical trend, with a saturation value of 1 being reached earlier for more massive hyperon species. This phenomenon underscores a preference for almost symmetric antihyperons-hyperons production in the case of multistrange hyperons, where the former, i.e., the $\Lambda$, contains only two light valence quarks, while the $\Xi$ contains only one light valence quark, the likely remnants from the initial nucleons. In the case of $\mathrm{\Omega}$, all (strange) valence quarks are newly generated Andronic2006 ; Bashir2015 . The generally unsatisfactory quality of thermal fits to the multi-strange baryons data at SPS energies cannot solely be attributed to these cases, as we shall see below Andronic2006 . All these ratios increase sharply with respect to $\sqrt{s_{NN}}$ and then almost saturate at higher energies (above 200 GeV), reaching a value equal to 1.0 at LHC energy. This behavior shows that the production rates of anti-particles relative to particles continuously increase with increasing $\sqrt{s_{NN}}$ and become almost equal at LHC energy. The excellent agreement between our VDW-HRG model results and the experimental data validates the ability of the model to explain these particle ratios.
IV.2 OTHER PARTICLE RATIOS
@@
We extend our analysis to further test the above values of the chemical freeze-out parameters. We find that the $K^{-}/K^{+}$ ratio is well described with the same values of the model parameters used to obtain the best fitted curves for antibaryon to baryon ratios. In Fig. 2 the $K^{-}/K^{+}$ ratio data against $\sqrt{s_{NN}}$ is well fitted by the curve obtained by using the VDW-HRG model calculations with $\chi^{2}/\text{{dof}}$ = 1.07 Abelev2010 ; Abelev2009 ; Acharya2020 ; Adams2002S ; Abelev2013 ; Mischke2002 . The decline in the $K^{-}/K^{+}$ ratio as collision energy decreases is attributed to a rise in net baryon density, favoring the production of $K^{+}$ (having $u$ and $\bar{s}$ valence quarks) over $K^{-}$ (having $\bar{u}$ and $s$ valence quarks). As the collision energy increases, the $K^{-}/K^{+}$ ratio exhibits a systematic rise and approaches unity, indicative of the pair production processes for $K^{+}$ and $K^{-}$ at higher energies in almost equal proportion through the reaction channels involving baryons and antibaryons in an almost baryon symmetric matter (having $\mu_{B}\sim 0$). The associated production mechanism yields $K^{+}$ through reactions like $\pi+N\rightarrow Y+K^{+}$ and the production of $K^{-}$ takes place through $\pi+\bar{N}\rightarrow\bar{Y}+K^{-}$, where $N$ represents a nucleon and $Y$ denotes a singly strange hyperon. On the other hand, reactions like $N+\bar{N}\rightarrow K^{-}+K^{+}$ and $\pi^{+}+\pi^{-}\rightarrow K^{-}+K^{+}$ will produce $K^{+}$ and $K^{-}$ in equal proportion Das2014 .
Figure 2: $\sqrt{s_{NN}}$ dependence of $\pi^{-}/\pi^{+}$, $K^{-}/K^{+}$
and $p/\pi$
For the purpose of comparison and using the same set of parameters we have also shown the energy dependence of the $\pi^{-}/\pi^{+}$ ratio in the same figure where both ratios approach $\sim 1$ for large $\sqrt{s_{NN}}$, as the system tends to become (almost) electric charge and baryon symmetric due to nuclear transparency effect. The green curve shows the theoretical fit, which matches the experimental data Bratkovskaya2004 ; Adams2002S ; Abelev2013 ; Adamczyk2017B ; Adler2002 quite well with $\chi^{2}/\text{{dof}}=1.25$. It may be recalled here that in our calculation we have applied the criteria of conserved charge to baryon ratio in the experiments in order to fix the electric chemical potential $\mu_{Q}$. We find that at low beam energies, the $\pi^{-}/\pi^{+}$ ratios exhibit values larger than unity, possibly due to significant contributions to the positive charges from protons and other resonances, particularly the $\Delta$ baryons. Further, since protons (antiprotons) are the lightest baryons (antibaryons) hence most high-mass baryons (antibaryons) decay into protons (antiprotons). The $p/\pi$ ratio, therefore, can characterizes well the overall baryon (antibaryon) production relative to the pions which are copiously produced in URNNC. To understand this, in Fig. 2 we have shown the $p/\pi^{+}$ and $\bar{p}/\pi^{-}$ ratios energy dependence. The theoretical curve obtained for the same set of the parameters values appears to fit the data quite well with $\chi^{2}/\text{{dof}}=0.71$ and $3.09$, respectively Nandi2020 ; Aggarwal2011S ; Adamczyk2017B ; Abelev2010 . The relatively larger value of $\chi^{2}/\text{{dof}}$ for the $\bar{p}/\pi^{-}$ case results due to very small error bars in the experimental data.
@@
Figure 4: $\sqrt{s_{NN}}$ dependence of the strangeness imbalance factor $\xi_{s}$
Despite the success of the HRG Model in describing many particle ratios, some unresolved issues still remain while describing the yields of the strange (antistrange) particles relative to the non-strange ones. For nearly a decade, the $K/\pi$, $\Lambda/\pi^{-}$, $\Xi/\pi^{-}$, $\Omega/\pi^{-}$, $\Phi/\pi^{-}$ and $\Phi/K^{-}$ ratios were not accurately described Beutler2008 ; Tawfik2016 . Pions being the lightest hadrons are produced copiously and carry the bulk of the entropy of the system. The dependence of the $K/\pi$ on $\sqrt{s_{NN}}$ has always been a subject of interest as it is a crucial indicator of the strangeness content of the system relative to its entropy in heavy-ion collisions and has implications for the quark-hadron phase transition. The observed increase and subsequent saturation of the $K/\pi$ ratio with increasing beam energies align with established trends and theoretical expectations Andronic2006 . As already mentioned we have assumed simultaneous freeze-out for all particles in our calculation. The resulting fit to the experimental $K^{-}/\pi^{-}$ ratio’s dependence on $\sqrt{s_{NN}}$ is shown in Fig. 3 in blue color. It is obvious that the theoretical curve provides an unsatisfactory fit to the experimental data Adams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 . However, such large disagreements have been also reported earlier which is attributed to strangeness imbalance in the system Castorina2016 ; Munzinger2003 ; Becattini2005 ; Bugaev2016 .
Figure 5: Variation of $K^{+}/\pi^{+}$ with $\sqrt{s_{NN}}$
The theoretical curve is seen to highly over-predict the experimental data at higher collision energies ($\sqrt{s_{NN}}>10\,\text{GeV}$). To explain this disagreement and further analyze other strange to non-strange particle ratios, we introduce a strangeness imbalance factor $\xi_{s}$ which is defined at different energies (say for the $i^{th}$) as follows Tawfik2021H ; Biswas2020 :
@@
In order to make it convenient to apply this factor to other particle ratios as well for any given collision energy, we have used a parameterization method to represent the energy dependence of $\xi_{s}$. We have used the following ansatz to best represent the obtained $\sqrt{s_{NN}}$ dependent values of $\xi_{s}$:
@@
The best spline fit provides:
$\alpha=0.469\pm 0.016$, $\beta=0.432\pm 0.006$,
$\gamma=8.83\pm 0.364\,\text{GeV}$, and $\delta=6.71\pm 0.36$.
In Fig. 4, we have shown the energy dependence of $\xi_{s}$. Utilizing the ansatz of Eq. 39, we have redefined the strangeness imbalance corrected strange (antistrange) particle number densities in our calculation at different collision energies as
@@
The quantity $h$ represents the number of strange and/or antistrange quarks in the given hadron. For singly strange particles (i.e., kaons, antikaons, lambda, antilambda, sigma, antisigma) we have $h=1$; for doubly strange particles (i.e., cascades, anticascades) as well as $\phi$ meson we take $h=2$; and similarly for triply strange particles (i.e., omega, antiomega) we have $h=3$. In Fig. 3, the corrected theoretical curve for the $K^{-}/\pi^{-}$ ratio using the above ansatz in Eq. 39 is shown by the red curve, which has a very small $\chi^{2}/\text{{dof}}=0.51$, resulting from the correction procedure. Moreover, using the values of $\xi_{s}$ obtained from Eq. 39, we have also calculated the strangeness imbalance corrected $K^{+}/\pi^{+}$ ratio for different collision energies. In Fig. 5, the theoretical curve for the $K^{+}/\pi^{+}$ ratio is shown in blue, having value of $\chi^{2}/\text{{dof}}=0.15$. Interestingly, the theoretical curve also explains the horn structure at lower energies, $\sqrt{s_{NN}}\sim 7-8\,\text{GeV}$. The rapid variation of the $(K^{+}/\pi^{+})$ and $(K^{-}/\pi^{-})$ ratios at lower energies reflects the threshold for strangeness production affected by the rapid initial increase in the temperature with $\sqrt{s_{NN}}$. A steep drop in the $K^{+}/\pi^{+}$ near the horn may arise due to the rapidly falling baryon chemical potential, as baryon-rich matter favors $K^{+}$ production, as discussed above. Though the rapid rise in temperature and the subsequent drop in baryon chemical potential with increasing collision energy are significant factors, they alone cannot fully explain the horn structure within the VDW-HRG model without the presence of the strangeness imbalance factor $\xi_{s}$ which is critical in capturing the observed behavior. However, due to the strangeness imbalance effect the like-mass anti-hyperon to hyperon as well as antikaon to kaon ratios are not modified, as their abundances in the system are equally affected and in their theoretical calculations the correction factor $\xi_{s}$ cancels out.
Figure 6: Dependence of $\Lambda(\bar{\Lambda})/\pi^{-}$, $\Xi/\pi$, and $\Omega/\pi^{-}$ on $\sqrt{s_{NN}}$
The observation that the $K^{-}/\pi^{-}$ and $K^{+}/\pi^{+}$ ratios in Figs. 3 and 5 approach the same value (i.e., $\sim 0.15$) towards the highest RHIC and LHC energies is also supported by the limiting values ($\sim 1$) of the $K^{-}/K^{+}$ and $\pi^{-}/\pi^{+}$ ratios seen in Fig. 2 at the highest energies Andronic2009Horn .
In Fig. 6, we have shown the results of our calculations along with the experimental data for several other particle ratios, i.e., $\Lambda(\bar{\Lambda})/\pi^{-}$, $\Xi(\bar{\Xi})/\pi^{-}(\pi^{+})$ and $\Omega/\pi$ while in Fig. 7 the variation of the $\Phi/\pi^{-}$ and $\Phi/K^{-}$ ratios with $\sqrt{s_{NN}}$ is shown.
We find that a reasonable fit is obtained in all these cases, thus highlighting the need for consideration of the strangeness imbalance effect in the system. The experimental data are taken from the references Adam2020 ; Abelev2010 ; Adamczyk2017B ; Antii2003 ; RecentRF ; Alt2008 ; Aggarwal2011S ; Friese2004 ; E9172003 ; Nandi2020 . The theoretical fits to the data sets give $\chi^{2}/\text{{dof}}$ = 1.02 (1.89) for $\Lambda(\bar{\Lambda})/\pi^{-}$, 5.52 (5.27) for $\Xi(\bar{\Xi})/\pi^{-}(\pi^{+})$ and 1.12 for $\Omega^{-}/\pi^{-}$. The larger values for $\Xi(\bar{\Xi})/\pi^{-}(\pi^{+})$ arise due to very small error bars associated with the LHC experimental data at 2.76 TeV. The corresponding values for the $\Phi/\pi^{-}$ and $\Phi/K^{-}$ are 2.61, 0.61 respectively.
Figure 7: Variation of $\Phi/\pi^{-}$, and $\Phi/K^{-}$ with $\sqrt{s_{NN}}$.
We have analyzed the correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios in URNNC over a wide energy range from 5 GeV to 2.76 TeV. The correlation is best described by a power-law type relationship as $K^{-}/K^{+}=\left(\bar{p}/p\right)^{\nu}$. This can serve as validation for any thermal model as it indicates how kaon production relates to net baryon density. In Fig. 8, we have shown the experimental data points along the with curves obtained from different approaches. The blue curve shows the result of our calculations. The pink curve is obtained by considering the light valence quark compositions, thus providing $\nu\approx 0.33$ but fits the data somewhat poorly. For comparison, if we represent the theoretical blue curve by a power-law fit, this provides $\nu\approx 0.17$, which contrasts sharply with the light valence quark composition approach. However, if we obtain a direct fit to the data using the power law, we get $\nu\approx 0.21$, which is quite close to our theoretical results. This correlation is therefore well described by the VDW-HRG type approach in the framework of the grand canonical ensemble-based thermal model. Experimental data at different energies are taken from the references Bearden2003 ; Stiles2006 ; Das2014 ; Abelev2009 .
@@
IV.3 ENERGY DEPENDENCE OF THERMAL FREEZE-OUT PARAMETERS ($T$, $\mu_{B}$, $\mu_{s}$, $\mu_{Q}$)